A car negotiate a bend of radius 20m with an acceleration of 12m/s^2. What is the maximum speed the car can attain without skidding?
1
Expert's answer
2020-09-08T09:11:16-0400

According to Newton's second law, if 12 m/s2 is centripetal acceleration:


"ma=m\frac{v^2}{R},\\
v=\sqrt{aR}=\sqrt{12\cdot20}=15.5\text{ m/s}."